# Constraining anti-baryonic dark matter through correlated nucleon decay signatures

### v1

@@ III Spontaneous breaking of dark baryon number and generation of di-nucleon decay
where $\Lambda_{nn\rightarrow\pi\pi}$ is the effective field theory scale. Matching the effective operator in Eq. (21) with the one-loop process in Fig. (4) gives,
$\frac{1}{\Lambda_{nn\rightarrow\pi\pi}}=\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\to\pi}^{2}}\frac{(\lambda_{1}v_{S})^{2}(\lambda_{2}v_{\Delta})}{M_{s}^{2}}=\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\to\pi}^{2}}\frac{m_{\Phi}^{3}}{M_{s}^{2}}\ .$ (22)
Using the decay width for the process,
$\Gamma_{nn\pi\pi}\sim\frac{m_{n}^{3}}{\Lambda_{nn\rightarrow\pi\pi}^{2}}\ ,$ (23)
and the lifetime for the decay $nn\rightarrow\pi\pi$ given in Table. 1, the constraint on the scale for this process becomes,
[…]
Assuming $\lambda_{1}v_{S}=\lambda_{2}v_{\Delta}=m_{\Phi}=10\hskip 2.84544pt\mathsf{GeV}$ and $M_{s}=0.1\ \Lambda_{n\rightarrow\bar{\nu}}$, the limit on $\Lambda_{n\rightarrow\bar{\nu}}$ becomes
$\Lambda_{n\to\pi}>2.99\times 10^{8}\hskip 2.84544pt\mathsf{GeV}\ .$ (25)
The spontaneous breaking of the baryon/lepton number in the dark sector, hence, leads to an independent constraint on the nucleon decay scale given by Eq. (19) and Eq. (25). Correlating these results with the induced nucleon decay discussed in Eq. (13) can then provide an additional handle for constraining the properties of dark matter.
@@ IV Correlating induced nucleon decay with di-nucleon decay
In the previous section, we derived the di-nucleon processes generated at one-loop from the induced nucleon decay processes. Furthermore, we used the experimental constraints on the di-nucleon decay widths computed in Eq. (17) and Eq. (23) to determine the new physics energy scales for the nucleon decay operators in Eq. (19) and Eq. (25). Importantly, these results are independent of the dark matter density on Earth.
In contrast, the decay widths of induced nucleon processes depend on the dark matter flux at Earth. In this section, we use the di-nucleon process to better constrain the dark matter densities at the terrestrial experiments.
For $n\rightarrow\bar{\nu}+\text{invisible}$ decay, the thermal cross-section in Eq. (8) computed using the constraint on the new physics scale from Eq. (19) becomes,
@@
$=$ $4.73\times 10^{-16}\hskip 2.84544pt\mathsf{GeV}^{-2}\ ,$
where we have used $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$. Substituting this into Eq. (13), we obtain the decay width of the induced nucleon decay as,
$\Gamma_{n\rightarrow\bar{\nu}}$ $=$ $\frac{\rho_{DM}}{M_{DM}}\cdot 4.73\times 10^{-16}\ \text{GeV}^{-2}\ .$ (27)
Using the experimental constraint on the decay width for $n\rightarrow\bar{\nu}+\text{invisible}$ given in Table. (1), the local terrestrial density of dark matter can be inferred as, $\rho_{DM}=9.86\times 10^{-5}\hskip 2.84544pt\mathsf{GeV\cdot cm^{-3}}$ for a $10\hskip 2.84544pt\mathsf{GeV}$ dark matter candidate.
Alternatively, by treating $\Lambda_{n\rightarrow\bar{\nu}}$ as a variable, while fixing the dark matter mass at $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$, one can compute the dependence of the terrestrial density of dark matter $\rho_{DM}$ on the new physics scale $\Lambda_{n\rightarrow\bar{\nu}}$. This dependence is illustrated in Fig. (5). The red vertical line in the plot indicates the lower bound on $\Lambda_{n\to\bar{\nu}}$ derived in Eq. (19). The region above the curve is excluded by the nucleon decay experiments, while the shaded region is disfavored by the limits derived from the $nn\rightarrow\bar{\nu}\bar{\nu}$ process.
Figure 5: The plot shows the local terrestrial dark matter density $\rho_{DM}$ vs the new physics scale $\Lambda_{n\rightarrow\nu}$ with the red vertical line indicating the limit from the $nn\rightarrow\bar{\nu}\bar{\nu}$ process. The blue curve represents the parameter space that satisfies the current limit for $n\to\bar{\nu}+\text{invisible}$, assuming a dark matter mass $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$  Huang:2013xfa . The parameter space above the curve is disallowed by $n\to\bar{\nu}+\text{invisible}$, while the shaded region is disallowed by $nn\rightarrow\bar{\nu}\bar{\nu}$.
For the scenario involving baryon-violating decay via $n\rightarrow\pi^{0}$, the thermal cross-section can be calculated using Eq. (12) and Eq. (25) as,
@@
$=$ $1.14\times 10^{-19}\hskip 2.84544pt\mathsf{GeV}^{-2}\ .$
@@
Using this result, the second expression in Eq. (13) for the induced nucleon decay width becomes,
$\Gamma_{n\rightarrow\pi}$ $=$ $\frac{\rho_{DM}}{M_{DM}}\cdot 1.14\times 10^{-19}\hskip 2.84544pt\mathsf{GeV}^{-2}\ .$ (29)
@@
Given the experimental value for the decay width of $n\rightarrow\pi^{0}$ provided in Table. (1), the corresponding local dark matter density can be computed to be $\rho_{DM}=2.16\times 10^{-4}\hskip 2.84544pt\mathsf{GeV/cm^{3}}$ for a $10\hskip 2.84544pt\mathsf{GeV}$ dark matter candidate.
As in the case of $n\to\bar{\nu}$, fixing $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$ and using the value of $\Gamma_{n\rightarrow\pi}$ from Table. (1), we can plot terrestrial $\rho_{DM}$ as a function of $\Lambda_{n\rightarrow\pi}$, as shown in fig. (5). The region above the curve is excluded by nucleon decay constraint, while the limits from the $nn\rightarrow\pi\pi$ process disfavor all values of $\Lambda_{n\rightarrow\pi}$ below $2.99\times 10^{8}\hskip 2.84544pt\mathsf{GeV}$, indicated by the vertical red line.
@@
Figure 6: The plot shows the local dark matter density $\rho_{DM}$ vs the $\Lambda_{n\rightarrow\pi}$, with the red vertical line indicating the limit from the $nn\rightarrow\pi^{0}\pi^{0}$ process. The blue curve represents the parameter space that satisfies the current limit for $n\to\pi^{0}+\text{invisible}$, assuming a dark matter mass $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$ Davoudiasl:2010am ; Davoudiasl:2011fj . The region above the blue curve is disallowed by $n\to\pi^{0}+\text{invisible}$, while the shaded region is disallowed by $nn\rightarrow\pi^{0}\pi^{0}$.
@@
Therefore, for $B-L$ preserving $n\rightarrow\bar{\nu}$ and $B-L$ violating $n\rightarrow\pi^{0}$ processes, the anti-baryonic dark matter density at Earth is summarized in Table. (4), assuming a dark matter mass of $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$. As is evident, such dark matter contributes only to a fraction of the total dark matter flux at Earth.
  Models    Decay Reaction   $\sigma v$   $\rho_{DM}$
$\hskip 5.69046pt(\mathsf{GeV}^{-2})\hskip 5.69046pt$ $\hskip 5.69046pt(\mathsf{GeV\cdot cm^{-3}})\hskip 5.69046pt$
Case-1 $n\rightarrow\bar{\nu}$ $\hskip 5.69046pt4.73\times 10^{-16}\hskip 5.69046pt$ $\hskip 5.69046pt9.86\times 10^{-5}\hskip 5.69046pt$
Case-2 $n\rightarrow\pi^{0}$ $\hskip 5.69046pt1.14\times 10^{-19}\hskip 5.69046pt$ $\hskip 5.69046pt2.16\times 10^{-4}\hskip 5.69046pt$
@@
Table 4: Benchmark points on the constraint on local dark matter density ($\rho_{DM}$) upon simultaneously satisfying the induced nucleon and di-nucleons decays as shown in Fig. (5) and Fig. (6) for dark matter mass $M_{DM}=10\hskip 2.84544pt\mathsf{GeV}$ Davoudiasl:2010am ; Davoudiasl:2011fj ; Huang:2013xfa .
@@
On the other hand, if we assume that the anti-baryon number charged dark matter constitutes the entirety of the local terrestrial dark matter flux ($\rho_{DM}=0.3\hskip 2.84544pt\mathsf{GeV}/\mathsf{cm^{3}}$ ), then the constraint from $n\to\bar{\nu}+\text{invisible}$ and $nn\to\bar{\nu}\bar{\nu}$ bounds the dark matter mass to $\leq 52\hskip 2.84544pt\mathsf{GeV}$ as shown by the intersection of the black curve in the Fig. (7) with the $\rho_{critical}$. Meanwhile, as indicated by the blue curve in the same figure, the constraint from $n\to\pi^{0}+\text{invisible}$ and $nn\to\pi^{0}\pi^{0}$ restricts the dark matter mass to $\leq 80\hskip 2.84544pt\mathsf{GeV}$. These results are summarized in the Table. (5).
@@
Figure 7: The plot shows the terrestrial dark matter density $\rho_{DM}$ as a function of the dark matter mass $M_{DM}$. The black (blue) curves represents the regions of parameter space that simultaneously satisfies the limits from $n\to\bar{\nu}+\text{invisible}$ and $nn\rightarrow\bar{\nu}\bar{\nu}$ ($n\to\pi^{0}+\text{invisible}$ and $nn\rightarrow\pi^{0}\pi^{0}$). The red horizontal line denotes to the total dark matter flux on Earth. The region to the left of each curve corresponds to new physics scales not yet probed by current di-nucleon searches. All regions to the right of the black/blue curve are excluded from the terrestrial nucleon decay experiments, while the region above the red horizontal line is excluded from dark matter flux constraints. Therefore, only the region below the curves and the red horizontal line is allowed.
@@
Case-1 $n\rightarrow\bar{\nu}$ $\hskip 5.69046pt9.9\times 10^{-19}\hskip 5.69046pt$ $\hskip 5.69046pt\leq 52\hskip 5.69046pt$
Case-2 $n\rightarrow\pi^{0}$ $\hskip 5.69046pt6.1\times 10^{-22}\hskip 5.69046pt$ $\hskip 5.69046pt\leq 80\hskip 5.69046pt$

### v2

@@ III Spontaneous breaking of dark baryon number and generation of di-nucleon decay
where $\Lambda_{nn\to\pi\pi}$ denotes the effective new-physics scale associated with the process. Matching the effective operator in Eq. (21) with the one-loop process in Fig. (4) we obtain the relation,
$\frac{1}{\Lambda_{nn\rightarrow\pi\pi}}$ $=$ $\frac{1}{16\pi^{2}}\frac{1}{\Lambda_{n\to\pi}^{2}}\frac{1}{3}\Big(\frac{\lambda_{1}\lambda_{2}v_{S}^{2}m_{\phi}}{M_{DM_{1}}M_{DM_{2}}}\Big)$ (22)
$=$ $\frac{1}{16\pi^{2}}\frac{M_{DM}}{\Lambda_{n\to\pi}^{2}}\frac{1}{3}R_{\pi\pi}\ ,$
where $M_{DM_{1}}^{2}=\lambda_{1}v_{S}m_{\phi}+m_{\phi}^{2}$ and $M_{DM_{2}}=\lambda_{2}v_{S}+m_{\psi}$ are the masses of the dark matter candidates. In the second line of the above equation, we have assumed $m_{\phi}\simeq m_{\psi}$ which allows us to identify $M_{DM_{1}}=M_{DM_{2}}=M_{DM}$. Furthermore, the dimensionless parameter,
$R_{\pi\pi}=\frac{\lambda_{1}\lambda_{2}v_{S}^{2}}{M_{DM}^{2}}\ll 1\ ,$
like before, incorporates the suppression associated with the spontaneous breaking of the baryon symmetry.
The decay width corresponding to the di-nucleon decay channel $nn\to\pi^{0}\pi^{0}$ becomes,
$\Gamma_{nn\pi\pi}\sim\frac{m_{n}^{3}}{\Lambda_{nn\rightarrow\pi\pi}^{2}}\ .$ (23)
Together with the experimental lower bound on the lifetime for the di-nucleon decay channel $nn\rightarrow\pi^{0}\pi^{0}$, given in Table. (1), the constraint on the scale for this process can be computed as,
$\frac{1}{\Lambda_{nn\rightarrow\pi\pi}^{2}}<6.11\times 10^{-63}\hskip 2.84544pt\mathsf{GeV}^{-2}\ .$ (24)
This bound, although highly suppressed, has important implications for scenarios where baryon number violating processes are mediated through the dark sector. In particular, assuming a benchmark dark matter mass $M_{DM}=2\ \mathsf{GeV}$ and $R_{\pi\pi}\sim 10^{-5}$ the independent lower bound on the scale of the induced neutron–pion decay operator, $\Lambda_{n\rightarrow\bar{\nu}}$ becomes,
$\Lambda_{n\to\pi}>1.8\times 10^{9}\hskip 2.84544pt\mathsf{GeV}\ .$ (25)
The spontaneous breaking of baryon and/or lepton number in the dark sector thus gives rise to an independent constraint on the scale of the induced nucleon decay operators. In particular, the bounds derived from the di-nucleon decay channels serve as complementary probes to those obtained from single nucleon decay processes. Correlating the limits in Eq.(19) and Eq.(25) with the decay width for induced nucleon decay in Eq.(13) provides a nontrivial cross-check on the parameter space of the model. Such a correlation not only constrains the effective scale of baryon number violation, but also offers an additional handle for restricting the local density ($\rho_{DM}$), masses ($M_{DM}$) and the strength of the spontaneous symmetry breaking ($\lambda_{i}$) of the dark sector fields.
@@ IV Correlating induced nucleon decay with di-nucleon decay
In the previous section, we derived the di-nucleon processes generated at one-loop from the induced nucleon decay operators. Furthermore, by employing the experimental constraints on the di-nucleon decay widths in Eq. (17) and Eq. (23), we extracted the corresponding limits on the new physics scales for the nucleon decay operators, as shown in Eq. (19) and Eq. (25).
This implies that while the nucleon decay constraints are sensitive to local density, the di-nucleon processes provide a model independent benchmark for the effective scales of baryon number violation. Consequently, the interplay between these two sets of observables enables us to translate the bounds from di-nucleon processes into nontrivial constraints on the local dark matter abundance relevant for laboratory experiments. In the following, we quantify this relation explicitly and demonstrate how the di-nucleon bounds sharpen the consistency requirements on induced nucleon decay signatures.
For $n\rightarrow\bar{\nu}+\text{invisible}$ decay, the thermal cross-section in Eq. (8) can be re-expressed using the bound on the new physics scale $\Lambda_{n\to\bar{\nu}}$ obtained in Eq.(19). Substituting this constraint, we get,
@@
$=$ $1.8\times 10^{-17}\ \mathsf{GeV}^{-2}\ ,$
where we have used $M_{DM}=2\ \mathsf{GeV}$. Substituting this into Eq. (13), the decay width of the induced nucleon process becomes,
$\Gamma_{n\rightarrow\bar{\nu}}$ $=$ $\frac{\rho_{DM}}{M_{DM}}\cdot 1.8\times 10^{-17}\ \text{GeV}^{-2}\ .$ (27)
Comparing with the experimental upper bound on the decay width for $n\rightarrow\bar{\nu}+\text{invisible}$, given in Table. (1), allows us to directly infer the local terrestrial density of dark matter. For a dark matter of mass $M_{DM}=2\ \mathsf{GeV}$, we find, $\rho_{DM}=0.5\times 10^{-3}\mathsf{GeV\cdot cm^{-3}}$, which lies well below the canonical astrophysical estimate $\rho_{DM}=0.3\ \mathsf{GeV\cdot cm^{-3}}$.
Alternatively, instead of fixing the scale $\Lambda_{n\rightarrow\bar{\nu}}$, one can treat it as a variable and determine how the terrestrial dark matter density $\rho_{DM}$ varies as a function of $\Lambda_{n\rightarrow\bar{\nu}}$. The resulting dependence is shown in Fig.(5). The red vertical line denotes the lower bound on $\Lambda_{n\to\bar{\nu}}$ obtained previously in Eq. (19). The region lying above the curve corresponds to parameter space excluded by existing nucleon decay searches, since it would predict a decay rate in excess of the experimental bounds. In addition, the shaded band indicates the portion of parameter space already disfavored by the complementary constraints arising from the di-nucleon decay channel $nn\rightarrow\bar{\nu}\bar{\nu}$.
Figure 5: The plot shows the local terrestrial dark matter density $\rho_{DM}$ vs the new physics scale $\Lambda_{n\rightarrow\nu}$ with the red vertical line indicating the limit from the $nn\rightarrow\bar{\nu}\bar{\nu}$ process. The blue curve represents the parameter space that satisfies the current limit for $n\to\bar{\nu}+\text{invisible}$, assuming a dark matter mass $M_{DM}=2\ \mathsf{GeV}$ Huang:2013xfa and $R_{\bar{\nu}\bar{\nu}}\sim 10^{-7}$. The parameter space above the curve is disallowed by $n\to\bar{\nu}+\text{invisible}$, while the shaded region is disallowed by $nn\rightarrow\bar{\nu}\bar{\nu}$.
For the case of baryon number violating decay via $n\rightarrow\pi^{0}$, the thermal cross-section can be computed using Eq. (12) together with the bound on the new physics sale Eq. (25) as,
@@
$=$ $2.5\times 10^{-20}\hskip 2.84544pt\mathsf{GeV}^{-2}\ .$
@@
Substituting this into the second expression in Eq. (13), the induced nucleon decay width becomes,
$\Gamma_{n\rightarrow\pi}$ $=$ $\frac{\rho_{DM}}{M_{DM}}\cdot 2.5\times 10^{-20}\hskip 2.84544pt\mathsf{GeV}^{-2}\ .$ (29)
@@
Using the experimental upper limit on the partial width for $n\rightarrow\pi^{0}$, listed in Table. (1), the corresponding local dark matter density can be inferred to be $\rho_{DM}=0.2\times 10^{-3}\ \mathsf{GeV/cm^{3}}$ for a dark matter candidate with mass $2\ \mathsf{GeV}$.
As before, treating $\Lambda_{n\to\pi}$ as a variable, Fig. (5) illustrates the terrestrial $\rho_{DM}$ as a function of the induced nucleon decay scale. The vertical red line denotes the lower bound on $\Lambda_{n\to\pi}$ obtained in Eq.(25). Thus, like in the previous case, the overlap between the nucleon-decay and di-nucleon constraints provides a stringent combined limit on baryon number violating dark sector interactions.
@@
Figure 6: The plot shows the local dark matter density $\rho_{DM}$ vs the $\Lambda_{n\rightarrow\pi}$, with the red vertical line indicating the limit from the $nn\rightarrow\pi^{0}\pi^{0}$ process. The blue curve represents the parameter space that satisfies the current limit for $n\to\pi^{0}+\text{invisible}$, assuming a dark matter mass $M_{DM}=2\ \mathsf{GeV}$ Davoudiasl:2010am ; Davoudiasl:2011fj and $R_{\pi\pi}\sim 10^{-5}$. The region above the blue curve is disallowed by $n\to\pi^{0}+\text{invisible}$, while the shaded region is disallowed by $nn\rightarrow\pi^{0}\pi^{0}$.
@@
Therefore, for the BLB-L preserving process $n\rightarrow\bar{\nu}$ and the BLB-L violating process $n\rightarrow\pi^{0}$, the corresponding anti-baryonic dark matter density are summarized in Table. (5), assuming a benchmark dark matter mass of $M_{DM}=2\ \mathsf{GeV}$.
 Models   Decay Reaction $\Lambda_{n}$ $\sigma v$ $\rho_{DM}$
$\hskip 5.69046pt(\mathsf{GeV})\hskip 5.69046pt$ $\hskip 5.69046pt(\mathsf{GeV}^{-2})\hskip 5.69046pt$ $\hskip 5.69046pt(\mathsf{GeV\cdot cm^{-3}})\hskip 5.69046pt$
Case-1 $n\rightarrow\bar{\nu}$ $4.3\times 10^{7}$ $\hskip 5.69046pt1.8\times 10^{-17}\hskip 5.69046pt$ $\hskip 5.69046pt0.5\times 10^{-3}\hskip 5.69046pt$
Case-2 $n\rightarrow\pi^{0}$ $2.2\times 10^{9}$ $\hskip 5.69046pt2.5\times 10^{-20}\hskip 5.69046pt$ $\hskip 5.69046pt0.3\times 10^{-3}\hskip 5.69046pt$
@@
Table 5: Benchmark points on the constraint on local dark matter density ($\rho_{DM}$) upon simultaneously satisfying the induced nucleon and di-nucleons decays as shown in Fig. (5) and Fig. (6) for dark matter mass $M_{DM}=2\ \mathsf{GeV}$ Davoudiasl:2010am ; Davoudiasl:2011fj ; Huang:2013xfa , assuming $R_{\bar{\nu}\bar{\nu}}\sim 10^{-7}$ and $R_{\pi\pi}\sim 10^{-5}$ respectively.
It is thus evident that the contribution from these anti-baryonic dark matter candidates constitutes only a small fraction of the total dark matter flux at Earth. While the nucleon decay constraints yield valuable insights into the scale of baryon-number violating operators, the parameter space consistent with terrestrial nucleon decay bounds leads to suppressed dark matter densities that cannot account for the entire local abundance.
@@
On the other hand, if we assume that the anti-baryon number charged dark matter constitutes the entirety of the local terrestrial dark matter flux ($\rho_{DM}=0.3\ \mathsf{GeV/cm^{3}}$), then the combined constraints from the processes $n\to\bar{\nu}+\text{invisible}$ and $nn\to\nu\nu$ bounds the dark matter mass to $\leq 14\ \mathsf{GeV}$ as shown by the intersection of the black curve in the Fig. (7) with the critical density $\rho_{critical}$. Meanwhile, as indicated by the blue curve in the same figure, the constraint from $n\to\pi^{0}+\text{invisible}$ and $nn\to\pi^{0}\pi^{0}$ restricts the dark matter mass to $\leq 20\ \mathsf{GeV}$. These results are summarized in the Table. (6).
@@
Figure 7: The plot shows the terrestrial dark matter density $\rho_{DM}$ as a function of the dark matter mass $M_{DM}$. The black (blue) curves represents the regions of parameter space that simultaneously satisfies the limits from $n\to\bar{\nu}+\text{invisible}$ and $nn\rightarrow\bar{\nu}\bar{\nu}$ ($n\to\pi^{0}+\text{invisible}$ and $nn\rightarrow\pi\pi$). The red horizontal line denotes the maximum local dark matter density on Earth.
@@
Case-1 $n\rightarrow\bar{\nu}$ $\hskip 5.69046pt2\times 10^{-19}\hskip 5.69046pt$ $\hskip 5.69046pt15\hskip 5.69046pt$
Case-2 $n\rightarrow\pi^{0}$ $\hskip 5.69046pt61.4\times 10^{-22}\hskip 5.69046pt$ $\hskip 5.69046pt20\hskip 5.69046pt$
@@
Moreover, the dependence of the local density of dark matter on $R_{\bar{\nu}\bar{\nu}}$ and $R_{\pi\pi}$ can be seen in Table.(7). It shows the lower limit on the local densities for dark matter of mass $M_{DM}=2\mathrm{GeV}$.
$R_{\bar{\nu}\bar{\nu}}$ $\Lambda_{n}$ $\rho_{DM}$
$\hskip 5.69046pt(\mathsf{GeV})\hskip 5.69046pt$ $\hskip 5.69046pt(\mathsf{GeV/cm^{3}})\hskip 5.69046pt$
$10^{-5}$ $\hskip 5.69046pt8.5\times 10^{8}\hskip 5.69046pt$ $\hskip 5.69046pt0.2\hskip 5.69046pt$
$10^{-6}$ $\hskip 5.69046pt8.5\times 10^{7}\hskip 5.69046pt$ $\hskip 5.69046pt0.002\hskip 5.69046pt$
$10^{-7}$ $\hskip 5.69046pt8.5\times 10^{6}\hskip 5.69046pt$ $\hskip 5.69046pt0.2\times 10^{-4}\hskip 5.69046pt$
$R_{\pi\pi}$ $\Lambda_{n}$ $\rho_{DM}$
$\hskip 5.69046pt(\mathsf{GeV})\hskip 5.69046pt$ $\hskip 5.69046pt(\mathsf{GeV/cm^{3}})\hskip 5.69046pt$
$3\times 10^{-4}$ $\hskip 5.69046pt7\times 10^{10}\hskip 5.69046pt$ $\hskip 5.69046pt0.27\hskip 5.69046pt$
$3\times 10^{-5}$ $\hskip 5.69046pt7\times 10^{9}\hskip 5.69046pt$ $\hskip 5.69046pt0.3\times 10^{-2}\hskip 5.69046pt$
$3\times 10^{-6}$ $\hskip 5.69046pt7\times 10^{8}\hskip 5.69046pt$ $\hskip 5.69046pt0.3\times 10^{-4}\hskip 5.69046pt$
Table 7: The lower limit on local density of the anti-baryonic dark matter ($M_{DM}=2$) for different $R_{\bar{\nu}\bar{\nu}}$ and $R_{\pi\pi}$.